Link to report from morning meeting
|Monday 1 October
- 03:13 beam dump by operation, natural end of fill !!!
- injection delayed, undulator problems
- fault on inj kicker MKI8_D (electronics)
|Sunday 20 September
- Injection bem 1 delayed, TI2 pc down
- 03:40 injection beam 1
- 05:22 Stable beams fill 3113, 1374b,
Luminosity ~ 5.6E33 cm-2s-1.
- 13:25 Beam dump, PC trip S67.
- 15:00 Filling again, target 1.5E11 p/b.
- 16:24 Stable beams fill 3114, 1374b,
Luminosity ~ 7E33 cm-2s-1.
|Saturday 29 September
- 00:00 Access for LHCb, CMS, then vacuum for PLC
- 03:30 End of access, precycle.
- 04:30 Injection - made it to 1374 bunches.
- 05:40 Stable beams fill 3110, 1374b, Luminosity
~ 5.3E33 cm-2s-1.
- 07:10 Beam dump, electrical glitch affecting S78 and
S81 due to UPS problem.
- UPS problem - Supplier intervention.
Thermal cycle on Q5 beam-screen. Access for LHCb, ATLA,
- 18:00 Preparing to inject.
- 19:20 OFC crash at flat top. Unfortunately this means
dumping the beams.
- 20:00 Short access for FBCT.
- 22:30 Injecting.
Friday 28 September
- 00:45 injection for physics
- 02:00 beam dump at ramp: T-alarm (fake) at collimators
- 06:30 injectio for physics
- 08:00 Stable beams fill 3108, 840b, Luminosity ~
- 09:40 Beam dump, trip of RCBXH3.R1.
- 10:30 Lost cryo S78.
- Access for LHCb and CMS in the shadow of the cryo
- 15:30 Cryo S78 back. Precycle.
- 17:00 Filling, managed to get up to 726b of 1.4E11.
- 18:40 Stable beams fill 3109, 726b, Luminosity
- 21:40 ATLAS VdM scans.
- 23:11 Beam dump, lost cryo S78 due to vacuum
Thursday 27 September
- scrubbing run for MKI8_D vacuum, injection of up to 5*144 bunches
- 09:00 injection for 2*456 fill
- 10:30 Stable beams fill 3102, 452b, Luminosity
- 11:45 beam dump: Q9.l2: water leak
- 18:30 machine closed, pre-cycle
- 20:00 injection delayed: 40MHz PS
|Wednesday 26 September
- 02:00 Cryo Pt back. RF problem on M1B1.
- 04:00 Cycle for loss maps with RPs.
- 06:00 Loss maps finished.
- 08:15 Injection for 78b full. BQM fixed.
- MKI8D vacuum issue. Scrubbing with 36b at injection.
- 12:30 Injection of 78b, ramp, squeeze.
- 14:00 Stable beams fill 3087, 78b, Luminosity ~3.1E32 cm-2s-1.
- 15:30 Moving to adjust mode for EoF studies.
- 17:45 Beam dump.
- Preparing for beam scrubbing at injection for MKI8.
|Tuesday 25 September
- 01:20 Cycle for loss maps. Loss on COD interlock
- 03:30 Next attempt for loss maps, 3 bunches /beam. Did
not find optimum in LHCb.
- 05:00 Loss map betatron and synch dump with RP in beam.
- 07:00 New cycle with 3 bunches / beam.
- 10:30 Found all IPs. IP8 based on last trims over past
- 12:00 Loss map in H plane and off-momentum loss map +500
Hz with RPs in beam.
- 14:00 Preparing for access in IR4 (BQM, BRST, BBQ/Headtail).
- 17:00 SIG transformer switching. ATLAS toroids trip,
filters in Pt2, 4 and 6.
- 17:30 Switch on filters Pt2. Cryo Pt2 off, some
circuits, ALICE solenoid and dipole.
Monday 24 September
- Cryo back at 1:00
- Powering tests R78/R81
- 05:20 Beam in
- 08:00 Pilot ramp & squeeze, collide. 50 Hz noise studies, active
filter tests on MBs and dipoles.
- 12:30 Dump.
- 14:00 Injection check of TCLIB in Pt2.
- 14:30 Q20 injection setup.
- 17:00 Problems with high losses at start of TI2/B1. Beam not
correctly positioned in the extraction channel. Corrected.
- 20:00 Resume Q20 injection, up to 144b with B1. Some issues with
satellites in the PS.
- 23:20 MKI8D flashover.